# Strong Gravitational Lensing Effects of the Rotating Short-Haired Black Hole and Constraints from EHT Observations

### v2

@@ Abstract
For the short hairs that have a significant impact only near the event horizon, studying their strong gravitational lensing effects is of great significance for revealing the properties of these hairs. In this study, we systematically investigated the strong gravitational lensing effects in the rotating short-haired black hole and constrained its hair parameter $Q_{m}$. Specifically, $Q_{m}$ causes the event horizon radius, photon - orbit radius, and impact parameter to be lower than those of the Kerr black hole. Regarding the lensing coefficients $\bar{a}$ and $\bar{b}$, as the spin parameter $a$ increases, $\bar{a}$ shows an increasing trend, while $\bar{b}$ shows a decreasing trend. In the observational simulations of M87* and Sgr A*, the angular position and angular distance of the relativistic image increase with the increase of $a$, while the magnification of the image shows an opposite trend. The existence of $Q_{m}$ only intensifies these trends. More importantly, the rotating short-haired black hole exhibits a significant difference in time delay compared to other black hole models. Especially in the simulation of M87*, the time delay deviation between the rotating short-haired black hole and the Kerr black hole, as well as the Kerr-Newman black hole, can reach dozens of hours. Through a comparative analysis with the observational data from the EHT, we effectively constrain the parameter space of the rotating short-haired black hole. The results indicate that this model has potential application prospects in explaining cosmic black hole phenomena and provides a possible theoretical basis for differentiating between different black hole models.
@@ I Introduction
The general theory of relativity predicts the existence of black holes(BH), but finding evidence of BH in the real universe is crucial for validating this theory. This has greatly stimulated the interest of physicists. In September 2015, the LIGO detector first captured gravitational wave signals produced by the merger of two BHs, providing strong evidence for the success of general relativity Abbott et al. (2016). In the real universe, most of the observed astrophysical BHs are rotating and fit well with the Kerr metric Teukolsky (2015); Tao et al. (2023). However, this does not rule out the possibility of a range of Kerr-like BHs Azreg-Aïnou (2014); Ghosh (2015). In fact, as Xu et al. stated in the literature Xu et al. (2021), regarding the Kerr BH as the astrophysical BH in the real universe is not entirely applicable. This is because BH in the real universe do not exist in isolation but are coupled with dark matter or other forms of fields, making the real universe more complex. Therefore, accurately describing astrophysical BH using standard Kerr BH may present certain challenges.
@@
If a type of Kerr-like BH solution that considers coupling with anisotropic matter could be found to describe astrophysical BH, it might be more meaningful. The short-haired BH solution provided in the literature Brown and Husain (1997) better fits the real universe environment because their short-hairy BH solution considers BH solutions coupled with anisotropic matter. Their static spherically symmetric BH solution violates the no-hair theorem. The no-hair theorem states Hawking (1972); Israel (1967) that within the framework of general relativity, the properties of a BH can be completely described by three parameters: mass $M$, charge $Q$, and angular momentum $J$. However, once Einstein’s gravity is coupled with other matter, BH solutions that violate the no-hair theorem may emerge, such as Lee et al. (2019); Greene et al. (1993); Bizon (1990); Ovalle et al. (2021). Therefore, studying the properties of short-hairy BH solutions Brown and Husain (1997) would be very interesting. However, the BH solution they provided is static and spherically symmetric. It would be more valuable to extend this solution to the rotating and axisymmetric case. This is because, on the one hand, most astrophysical BHs are rotating, so a rotating solution can better match astrophysical BHs. On the other hand, jet phenomena are common around rotating BHs, and observing these jets helps us study and understand the physical behavior near the event horizon, especially for this type of short hair that has a significant impact near the event horizon. Based on this objective, Tang and Xu in 2022 used the Newman-Janis (NJ) algorithm to extend the short-hairy BH to the case of rotating short-hairy BH, and they also studied the impact of short-hair parameter on the BH shadow Tang and Xu (2022). Studying other properties of the rotating short-hairy BH is also very meaningful. For example, exploring their physical behavior under strong gravitational lensing can help us further understand the characteristics of rotating short-hairy BH and provide a window to test the no-hair theorem.
@@
Gravitational lensing is an important tool for studying black holes (or other massive objects) and their surrounding environments. It has been widely applied not only in understanding the structure of spacetime (e.g., Liebes (1964); Mellier (1999); Guzik et al. (2010); Schmidt (2008)), but also plays a crucial role in the search for dark matter (see Massey et al. (2010); Vegetti et al. (2023); Diaz Rivero and Dvorkin (2020); Şengül and Dvorkin (2022); Fairbairn et al. (2023), etc.). Meanwhile, it is an effective means to test gravitational theories (see Zhao et al. (2024); Bekenstein and Sanders (1994); Islam et al. (2020); Kumar et al. (2020); Chen and Jing (2009), etc.). In a strong gravitational field, such as near a BH, when light rays approach the BH (especially near the photon sphere), the light rays will be strongly bent, and may even orbit the black hole one or more times, that is, the deflection angle of the light rays exceeds $2\pi$. In this case, shadows, photon rings, and relativistic images will appear Gralla et al. (2019); Bozza (2010); Bozza et al. (2001); Bozza (2003). Regarding strong gravitational lensing, Virbhadra and Ellis analyzed the strong gravitational effects caused by a Schwarzschild BH through numerical simulations Virbhadra and Ellis (2000). Subsequently, Bozza Bozza et al. (2001); Bozza (2002, 2010) and Tsukamoto Tsukamoto (2017), among others, generalized it to a general static spherically symmetric spacetime. Of course, other static spherically symmetric BHs have also been studied accordingly (e.g., Man and Cheng (2015); Wei et al. (2015); Eiroa and Sendra (2013, 2012); Eiroa (2006, 2005); Zhao and Xie (2017); Zhao et al. (2024)). In addition, rotating axisymmetric black holes have also received extensive attention, such as Bozza (2008a, 2010, 2003); Wei et al. (2012); Islam and Ghosh (2021); Islam et al. (2021); Hsiao et al. (2020).
@@
In fact, in 2019, the EHT successfully captured images of the supermassive BH $M87^{*}$ Akiyama et al. (2019a), and in 2022, it captured images of the BH at the center of the Milky Way, $SgrA^{*}$ Akiyama et al. (2022a). These breakthrough observations have triggered an upsurge of research on strong gravitational lensing. The EHT’s observations provide a new laboratory for studying the properties of black holes, such as the black hole’s shadow, the jet phenomenon of the accretion disk, the event horizon, and other properties that previously could only be theoretically calculated but can now be verified through observation. Naturally, studying the strong gravitational lensing effects around BH is also of great significance, as it can magnify and distort the images of background celestial bodies, providing a unique method to detect BHs and their surrounding material distribution. Therefore, studying strong gravitational lensing in BHs that are more akin to those in the real universe holds greater physical value. The environment considered for the rotating short-haired BH Tang and Xu (2022) is more similar to the real universe (anisotropic matter). Studying strong gravitational lensing in such BH can help us test the no-hair theorem, distinguish Kerr BH, and further understand the properties of short-hair near the event horizon. For this kind of hair, which has significant effects near the event horizon but is difficult to observe for distant observers, observing the behavior of light reaching near the event horizon will help us better understand the properties of this short-hair.
@@ II Rotating Short-Haired Black Hole
In the framework of classical general relativity, the no-hair theorem states that the properties of a BH can be fully described by just three parameters: mass $M$, charge $Q$, and angular momentum $J$ Hawking (1972); Israel (1967). However, research indicates that the no-hair theorem might not be universal, and numerous studies have sought counterexamples. In the 1990s, Piotr Bizon was the first to discover a BH solution with ”hair” through numerical analysis, challenging the absoluteness of the no-hair theorem Bizon (1990). Ovalle et al. constructed a spherically symmetric BH solution with hair via the gravitational decoupling method Ovalle et al. (2021), and this was later extended to the rotating case Contreras et al. (2021). Additionally, other researchers have proposed BH solutions with hair based on scalar fields or other theoretical frameworks, such as Caramês (2023); Bakopoulos and Nakas (2023); Zhang (2022); Brihaye and Ducobu (2016). In the literature Brown and Husain (1997), the authors obtained short-haired BH solutions by coupling Einstein’s gravity with anisotropic matter. These solutions can present de Sitter and Reissner-Nordström BHs in some cases, and in others, they yield short-haired BH solution Brown and Husain (1997). The corresponding metric is given by
@@
Considering that actual BHs in the universe usually have rotational characteristics, it is particularly important to generalize such short-haired BH solutions to the rotating case. Against this background, Tang and Xu generalized the static spherically symmetric short-haired BH metric to the rotating case through the Newman-Janis (NJ) algorithm, obtained the rotating short-haired BH solution, and analyzed the influence of the short- hair parameter on the BH shadow Tang and Xu (2022). This generalization makes it possible to study the physical properties of short-haired BHs in a rotating background. Since the spacetime of a rotating short-hairedBH is closer to that of the real universe, the matter distribution within it is non-vacuum and anisotropic. This model provides a basis for further exploring the properties of rotating short-haired BHs. The metric form of the rotating short-haired BH is Tang and Xu (2022)
@@
Here, $a$ denotes the black hole’s spin parameter, and $Q_{m}$ represents the hair strength parameter. As analyzed in the original literature Brown and Husain (1997), the short-haired BH satisfies the Weak Energy Condition (WEC). Thus, the value of the parameter $k$ must meet the requirement $2k-1\geq 0$, meaning $k\geq\frac{1}{2}$. Moreover, the trace of the energy momentum tensor $T=2\rho(k-1)>0$ (this holds when $k>1$), indicating that the short-haired BH does not violate the no ”short-hair” theorem Nunez et al. (1996); Acharya and Sarkar (2024). When $k=1$, the metric (2) simplifies to the classical Kerr-Newman BH. When $k>1$, the metric (2) depicts a short-haired BH. In the discussion of this paper, we will concentrate on analyzing the BHs corresponding to $k=1$ and $k=\frac{3}{2}$, with particular emphasis on the case of $k=\frac{3}{2}$. The reason is that as the value of $k$ increases, the influence of the hair strength parameter $Q_{m}$ gradually diminishes, and the properties of the short-haired BH gradually approximate those of the Kerr BH, making it difficult to distinguish between them. Therefore, choosing $k=\frac{3}{2}$ as the main focus of our study can more clearly expose the characteristics of the short-haired BH and its differences from classical BHs (such as the Kerr-Newman BH with $k=1$ and the Kerr BH with $Q_{m}=0$).
To facilitate subsequent analysis, we’ve performed a non-dimensionalization of the metric (2), using $2M$ as the unit to non-dimensionalize physical quantities (e.g., $r\rightarrow 2Mx$). In this framework, for the parameter models under discussion, the event horizon of the rotating short-haired BH can be determined by the condition $g^{rr}=0$, which is equivalent to $\Delta=0$. As depicted in Figure 1, the distribution characteristics of the event horizons of rotating short-haired BHs under various conditions are presented. In the figure, the black curve represents the event horizon of the Kerr BH. It can be observed that for different values of the parameter $k$, the presence of the hair parameter $Q_{m}$ significantly impacts the position of the event horizon. Compared to the Kerr BH, the event horizon of the rotating short-haired BH is always smaller. This indicates that the introduction of the hair parameter $Q_{m}$ not only alters the spacetime structure of the BH but also enables in-depth study of the physical properties of the hair parameter within a strong gravitational field environment. Consequently, it offers a new theoretical perspective for understanding the characteristics of short-haired BHs.
[…]
III Strong Gravitational Lensing in the Spacetime of the Rotating Short-hairy Black Hole
@@
The condition for an unstable Photon orbit is given by Harko et al. (2009)
@@
Generally, the above equation has multiple solutions. Only the largest one is outside the event horizon radius, so it is defined as the radius of the unstable photon orbit. This can be clearly seen from the graph of the effective potential. As shown in Figure 2, when the impact parameter reaches the critical value (the green dashed line), the radius of the corresponding unstable photon orbit is marked as $x=x_{m}$. In the following discussion, we’ll use $x_{m}$ to represent the radius of the unstable photon orbit. As presented in Figure 3, because of the hair parameter $Q_{m}$, the radius $x_{m}$ of the unstable photon orbit differs significantly among various models. As the spin parameter increases, $x_{m}$ in different models gradually drops. Besides, the $x_{m}$ of the rotating short-haired BH is smaller than that of the Kerr BH (for the Kerr black hole, $Q_{m}=0$).
Figure 2: The effective potential functions for specific values under different models and the positions of the unstable photon orbits. The intersection point of the green dashed line and the black dashed line represents the position of the unstable photon orbit. Here, we take $Q_{m}=0.3$ and $a=0.2$. From left to right, they correspond to the black - hole models with $k=1$ and $k=1.5$ respectively. We plot the effective potential under different hair parameters.
Figure 3: The variation of the radius of the unstable photon sphere in different BH models with the spin parameter $a$ and the short-hair parameter $Q_{m}$. From left to right, it corresponds to the Kerr-Newman BH with $k=1$ and the rotating short-haired BH with $k=1.5$ respectively.
@@
Figure 4: The variation of the impact parameter with the spin parameter $a$ and the short hair parameter $Q_{m}$ in two BH models. From left to right, they correspond to the Kerr-Newman BH with $k=1$ and the rotating short-haired BH with $k=1.5$ respectively.
In the strong field limit, the deflection angle $\alpha_{D}$ can be given by Bozza (2002, 2003)
@@
To handle the integral (23), we draw on the approaches presented in the references Bozza (2002); Tsukamoto (2017). Specifically, we expand the deflection angle in the vicinity of the unstable photon sphere. Additionally, we utilize an intermediate quantity re-defined by scholar Naoki Tsukamoto under the strong field limit to reconstruct the integral range. The form of this variable can be found in the definition given in Tsukamoto (2017), and its expression is
@@
This variable has been well applied in corresponding literature, such as Fu and Zhang (2022); Islam and Ghosh (2021). By using the above intermediate variable, equation (23) becomes
@@
According to the description of the strong - field limit, when light approaches the vicinity of the photon orbit radius $x_{m}$, the deflection angle of the light increases sharply. At this point, the deflection angle can be described by an analytical expansion as Bozza (2002, 2003); Tsukamoto (2017)
@@
Theoretically, substituting the corresponding expressions should yield an analytical expression for the deflection angle. However, due to the excessive complexity and length of the expression, we have instead plotted the relationships among the lensing coefficients, the short hair parameter, and the spin parameter. As shown in Figure 5, the deflection coefficient $\bar{a}$ increases with the increase of the spin parameter $a$, and the presence of the hair parameter makes this growth more rapid. When the hair parameter vanishes, i.e., $Q_{m}=0$, the rotating short-haired BH degenerates into a Kerr BH. The behavior of the deflection coefficient $\bar{b}$ is exactly opposite to that of the coefficient $\bar{a}$, and the presence of hair causes the corresponding coefficient to decrease more rapidly. When the parameters $a=0$ and $Q_{m}=0$, the short-haired BH degenerates into a Schwarzschild BH. At this point, the lensing coefficient $\bar{a}=1$ and $\bar{b}=-0.4002$, which is in good agreement with the lensing coefficient values of the Schwarzschild BH Bozza (2002)(see the intersections of the red contour lines and the abscissa in the left and right plots of Figure 5, as well as Table 1 ). When the parameters $a=0$, $k=1$, and $Q_{m}\neq 0$, the short-haired BH degenerates into a Reissner-Nordström (RN) BH. Our calculated results are in excellent agreement with those calculated by Bozza Bozza (2002) (Table 1).
In addition, in Figure 6, we have plotted the variation of the deflection angle of the short-haired BH with the impact parameter under different values of the hair parameter and the spin parameter. Evidently, as the impact parameter $u$ continuously decreases, different hair parameters always correspond to different divergence points (the points on the dashed lines in Figure 6 correspond to the values of the impact parameter at divergence). Among them, the deflection angle of the Kerr BH is greater than that of the rotating short-haired BH under the same impact parameter (the black dashed line in Figure 6 represents the case of the Kerr BH). Overall, the presence of the hair parameter causes the deflection angle to diverge at a smaller impact parameter. Meanwhile, the variation trends of both the lensing coefficients $\bar{a}$ and $\bar{b}$, as well as the deflection angle, are similar to those of the standard Kerr BH.

### v3

@@ Abstract
For the short hairs that have a significant impact only near the event horizon, studying their strong gravitational lensing effects is of great significance for revealing the properties of these hairs. In this study, we systematically investigated the strong gravitational lensing effects in the rotating short-haired black hole and constrained its hair parameter $Q_{m}$. Specifically, $Q_{m}$ causes the event horizon radius, photon - orbit radius, and impact parameter to be lower than those of the Kerr black hole. Regarding the lensing coefficients $\bar{a}$ and $\bar{b}$, as the spin parameter $a$ increases, $\bar{a}$ shows an increasing trend, while $\bar{b}$ shows a decreasing trend. In the observational simulations of M87* and Sgr A*, the angular position and angular separation of the relativistic image increase with the increase of $a$, while the magnification of the image shows an opposite trend. The existence of $Q_{m}$ only intensifies these trends, while parameter $k$ suppresses such tendencies. More importantly, the rotating short-haired black hole exhibits a significant difference in time delay compared to other black hole models. Especially in the simulation of M87*, the time delay deviation between the rotating short-haired black hole and the Kerr black hole, as well as the Kerr-Newman black hole, can reach dozens of hours. Through a comparative analysis with the observational data from the EHT, we effectively constrain the parameter space of the rotating short-haired black hole. The results indicate that this model has potential application prospects in explaining cosmic black hole phenomena and provides a possible theoretical basis for differentiating between different black hole models.
@@ I Introduction
The general theory of relativity predicts the existence of black holes(BH), but finding evidence of BH in the real universe is crucial for validating this theory. This has greatly stimulated the interest of physicists. In September 2015, the LIGO detector first captured gravitational wave signals produced by the merger of two BHs, providing strong evidence for the success of general relativity [1]. In the real universe, most of the observed astrophysical BHs are rotating and fit well with the Kerr metric [2, 3]. However, this does not rule out the possibility of a range of Kerr-like BHs [4, 5]. In fact, as Xu et al. stated in the literature [6], regarding the Kerr BH as the astrophysical BH in the real universe is not entirely applicable. This is because BH in the real universe do not exist in isolation but are coupled with dark matter or other forms of fields, making the real universe more complex. Therefore, accurately describing astrophysical BH using standard Kerr BH may present certain challenges.
@@
If a type of Kerr-like BH solution that considers coupling with anisotropic matter could be found to describe astrophysical BH, it might be more meaningful. The short-haired BH solution provided in the literature [7] better fits the real universe environment because their short-hairy BH solution considers BH solutions coupled with anisotropic matter. Their static spherically symmetric BH solution violates the no-hair theorem. The no-hair theorem states [8, 9] that within the framework of general relativity, the properties of a BH can be completely described by three parameters: mass $M$, charge $Q$, and angular momentum $J$. However, once Einstein’s gravity is coupled with other matter, BH solutions that violate the no-hair theorem may emerge, such as [10, 11, 12, 13]. Therefore, studying the properties of short-hairy BH solutions [7] would be very interesting. However, the BH solution they provided is static and spherically symmetric. It would be more valuable to extend this solution to the rotating and axisymmetric case. This is because, on the one hand, most astrophysical BHs are rotating, so a rotating solution can better match astrophysical BHs. On the other hand, jet phenomena are common around rotating BHs, and observing these jets helps us study and understand the physical behavior near the event horizon, especially for this type of short hair that has a significant impact near the event horizon. Based on this objective, Tang and Xu in 2022 used the Newman-Janis (NJ) algorithm to extend the short-hairy BH to the case of rotating short-hairy BH, and they also studied the impact of short-hair parameter on the BH shadow [14]. Studying other properties of the rotating short-hairy BH is also very meaningful. For example, exploring their physical behavior under strong gravitational lensing can help us further understand the characteristics of rotating short-hairy BH and provide a window to test the no-hair theorem.
@@
Gravitational lensing is an important tool for studying BHs (or other massive objects) and their surrounding environments. It has been widely applied not only in understanding the structure of spacetime (e.g., [15, 16, 17, 18]), but also plays a crucial role in the search for dark matter (see [19, 20, 21, 22, 23], etc.). Meanwhile, it is an effective means to test gravitational theories (see [24, 25, 26, 27, 28], etc.). In a strong gravitational field, such as near a BH, when light rays approach the BH (especially near the photon sphere), the light rays will be strongly bent, and may even orbit the BH one or more times, that is, the deflection angle of the light rays exceeds $2\pi$. In this case, shadows, photon rings, and relativistic images will appear [29, 30, 31, 32]. Regarding strong gravitational lensing, Virbhadra and Ellis analyzed the strong gravitational effects caused by a Schwarzschild BH through numerical simulations [33]. Subsequently, Bozza [31, 34, 30] and Tsukamoto [35], among others, generalized it to a general static spherically symmetric spacetime. Of course, other static spherically symmetric BHs have also been studied accordingly (e.g., [36, 37, 38, 39, 40, 41, 42, 24]). In addition, rotating axisymmetric BHs have also received extensive attention, such as [43, 30, 32, 44, 45, 46, 47].
@@
In fact, in 2019, the EHT successfully captured images of the supermassive BH $M87^{*}$ [48], and in 2022, it captured images of the BH at the center of the Milky Way, $SgrA^{*}$ [49]. These breakthrough observations have triggered an upsurge of research on strong gravitational lensing. The EHT’s observations provide a new laboratory for studying the properties of BHs, such as the black hole’s shadow, the jet phenomenon of the accretion disk, the event horizon, and other properties that previously could only be theoretically calculated but can now be verified through observation. Naturally, studying the strong gravitational lensing effects around BH is also of great significance, as it can magnify and distort the images of background celestial bodies, providing a unique method to detect BHs and their surrounding material distribution. Therefore, studying strong gravitational lensing in BHs that are more akin to those in the real universe holds greater physical value. The environment considered for the rotating short-haired BH [14] is more similar to the real universe (anisotropic matter). Studying strong gravitational lensing in such BH can help us test the no-hair theorem, distinguish Kerr BH, and further understand the properties of short-hair near the event horizon. For this kind of hair, which has significant effects near the event horizon but is difficult to observe for distant observers, observing the behavior of light reaching near the event horizon will help us better understand the properties of this short-hair.
@@ II Rotating Short-Haired Black Hole
In the framework of classical general relativity, the no-hair theorem states that the properties of a BH can be fully described by just three parameters: mass $M$, charge $Q$, and angular momentum $J$ [8, 9]. However, research indicates that the no-hair theorem might not be universal, and numerous studies have sought counterexamples. In the 1990s, Piotr Bizon was the first to discover a BH solution with ”hair” through numerical analysis, challenging the absoluteness of the no-hair theorem [12]. Ovalle et al. constructed a spherically symmetric BH solution with hair via the gravitational decoupling method [13], and this was later extended to the rotating case [50]. Additionally, other researchers have proposed BH solutions with hair based on scalar fields or other theoretical frameworks, such as [51, 52, 53, 54]. In the literature [7], the authors obtained short-haired BH solutions by coupling Einstein’s gravity with anisotropic matter. These solutions can present de Sitter and Reissner-Nordström BHs in some cases, and in others, they yield short-haired BH solution [7]. The corresponding metric is given by
@@
Considering that actual BHs in the universe usually have rotational characteristics, it is particularly important to generalize such short-haired BH solutions to the rotating case. Against this background, Tang and Xu generalized the static spherically symmetric short-haired BH metric to the rotating case through the Newman-Janis (NJ) algorithm, obtained the rotating short-haired BH solution, and analyzed the influence of the short- hair parameter on the BH shadow [14]. This generalization makes it possible to study the physical properties of short-haired BHs in a rotating background. Since the spacetime of a rotating short-haired BH is closer to that of the real universe, the matter distribution within it is non-vacuum and anisotropic. This model provides a basis for further exploring the properties of rotating short-haired BHs. The metric form of the rotating short-haired BH is [14]
@@
Here, $a$ denotes the black hole’s spin parameter, and $Q_{m}$ represents the hair strength parameter. As analyzed in the original literature [7], the short-haired BH satisfies the Weak Energy Condition (WEC). Thus, the value of the parameter $k$ must meet the requirement $2k-1\geq 0$, meaning $k\geq\frac{1}{2}$. Moreover, the trace of the energy momentum tensor $T=2\rho(k-1)>0$ (this holds when $k>1$), indicating that the short-haired BH does not violate the no ”short-hair” theorem [55, 56]. When $k=1$, the metric (2) simplifies to the classical Kerr-Newman BH. When $k>1$, the metric (2) depicts a short-haired BH. In the following discussion, we will focus on analyzing the short-haired BH models corresponding to parameters $k=1$, $3/2$, and $k=2$.
To facilitate subsequent analysis, we’ve performed a non-dimensionalization of the metric (2), using $2M$ as the unit to non-dimensionalize physical quantities (e.g., $r\rightarrow 2Mx$). In this framework, for the parameter models under discussion, the event horizon of the rotating short-haired BH can be determined by the condition $g^{rr}=0$, which is equivalent to $\Delta=0$. As shown in Figure 1, we illustrate the distribution characteristics of the event horizon for the rotating short-haired BH, where the black curve represents the event horizon of the Kerr BH. Evidently, the introduction of the hair parameter $Q_{m}$ significantly alters the positional distribution of the event horizon. For different parameters $k$, the analysis shows that as the value of $k$ increases, the event horizon radius gradually converges toward the Kerr BH event horizon. This indicates that in the large $k$ value limit, the characteristics of the rotating short-haired BH progressively approach those of the Kerr BH. Compared to the Kerr BH, the event horizon radius of the rotating short-haired BH is always smaller. This phenomenon suggests that the introduction of the hair parameter not only changes the spacetime structure of the BH but also provides the possibility to study the physical properties of the hair parameter in strong gravitational field environments, thereby offering a new theoretical perspective for understanding the characteristics of the rotating short-haired BH.
Figure 1: The event horizons and Cauchy horizons under different parameter conditions. The solid lines represent the event horizons $x_{+}$ , and the dashed lines represent the Cauchy horizons $x_{-}$.
III Strong Gravitational Lensing in the Spacetime of the Rotating Short-haired Black Hole
@@
The condition for an unstable Photon orbit is given by [57]
@@
Generally, the above equation has multiple solutions. Only the largest one is outside the event horizon radius, so it is defined as the radius of the unstable photon orbit. This can be clearly seen from the graph of the effective potential. As shown in Figure 2, when the impact parameter reaches the critical value (the green dashed line), the radius of the corresponding unstable photon orbit is marked as $x=x_{m}$. In the following discussion, we’ll use $x_{m}$ to represent the radius of the unstable photon orbit. As presented in Figure 3, because of the hair parameter $Q_{m}$, the radius $x_{m}$ of the unstable photon orbit differs significantly among various models. As the spin parameter increases, $x_{m}$ in different models gradually drops. Furthermore, under the same parameter conditions, the $x_{m}$ value of a rotating short-haired BH is smaller than the corresponding value for a Kerr BH ($Q_{m}=0$), and as the parameter $k$ increases, the $x_{m}$ value of the rotating short-haired BH gradually approaches the corresponding value of the Kerr BH. This indicates that the parameter $k$ has the effect of suppressing the hair parameter effect, and when the $k$ value is large, the behavior of the rotating short-haired BH will degenerate to that of the Kerr BH. The red dashed line in the figure represents the Kerr BH case.
Figure 2: The effective potential under different BH models, where the intersection point of the green dashed line and the black dashed line represents the position of unstable photon orbits. Here we take $Q_{m}=0.3$, $a=0.2$, and from left to right correspond to BH models with $k=1$, $1.5$, $2$.
Figure 3: The variation of unstable photon orbit radius with the spin parameter a and hair parameter under different BH models, from left to right corresponding to BH models with $k=1$, $1.5$, and $2$. The red dashed line in the figure represents the Kerr BH case.
@@
Figure 4: The variation of impact parameter with the spin parameter $a$ and hair parameter under different BH models, from left to right corresponding to BH models with $k=1$, $1.5$, and $2$. The red dashed line in the figure represents the Kerr BH case.
In the strong field limit, the deflection angle $\alpha_{D}$ can be given by [34, 32]
@@
To handle the integral (23), we draw on the approaches presented in the references [34, 35]. Specifically, we expand the deflection angle in the vicinity of the unstable photon orbit radius. Additionally, we utilize an intermediate quantity re-defined by scholar Naoki Tsukamoto under the strong field limit to reconstruct the integral range. The form of this variable can be found in the definition given in [35], and its expression is
@@
This variable has been well applied in corresponding literature, such as [58, 45]. By using the above intermediate variable, equation (23) becomes
@@
According to the description of the strong - field limit, when light approaches the vicinity of the photon orbit radius $x_{m}$, the deflection angle of the light increases sharply. At this point, the deflection angle can be described by an analytical expansion as [34, 32, 35]
@@
Theoretically, substituting the corresponding expressions should yield an analytical expression for the deflection angle. However, due to the excessive complexity and length of the expression, we have instead plotted the relationships among the lensing coefficients, the short hair parameter, and the spin parameter. As shown in Figure 5, the deflection coefficient $\bar{a}$ increases with the increase of the spin parameter $a$, and the presence of the hair parameter makes this growth more rapid. When the hair parameter vanishes, i.e., $Q_{m}=0$, the rotating short-haired BH degenerates into a Kerr BH. The behavior of the deflection coefficient $\bar{b}$ is exactly opposite to that of the coefficient $\bar{a}$, and the presence of hair causes the corresponding coefficient to decrease more rapidly. When the parameters $a=0$ and $Q_{m}=0$, the short-haired BH degenerates into a Schwarzschild BH. At this point, the lensing coefficient $\bar{a}=1$ and $\bar{b}=-0.4002$, which is in good agreement with the lensing coefficient values of the Schwarzschild BH [34](see the intersections of the red contour lines and the abscissa in the left and right plots of Figure 5, as well as Table 1 ). When the parameters $a=0$, $k=1$, and $Q_{m}\neq 0$, the short-haired BH degenerates into a Reissner-Nordström (RN) BH. Our calculated results are in excellent agreement with those calculated by Bozza [34] (Table 1). Furthermore, from the data in Table 1, it can be clearly observed that under the same conditions, as the parameter $k$ value increases, its corresponding coefficient values gradually approach the parameter values of the Kerr BH ($Q_{m}=0$). This indicates that the increase in parameter $k$ weakens the influence of the hair parameter, causing the properties of the rotating short-haired BH to gradually approach those of the Kerr BH.
In addition, in Figure 6, we have plotted the variation of the deflection angle of the short-haired BH with the impact parameter under different values of the hair parameter and the spin parameter. Evidently, as the impact parameter $u$ continuously decreases, different hair parameters always correspond to different divergence points (the points on the dashed lines in Figure 6 correspond to the values of the impact parameter at divergence). Among them, the deflection angle of the Kerr BH is greater than that of the rotating short-haired BH under the same impact parameter (the black dashed line in Figure 6 represents the case of the Kerr BH). In general, the introduction of the hair parameter leads to divergent characteristics in the deflection angle at smaller impact parameters. Notably, both the lensing coefficients $\bar{a}$ and $\bar{b}$, as well as the deflection angle $\alpha$, exhibit changes that show similarity to the standard Kerr BH. Further observation reveals that as the parameter $k$ increases, the trends of these physical quantities gradually converge with the behavior of the Kerr BH. Additional analysis indicates that when $k$ takes larger values, the physical characteristics of rotating short-haired BHs will tend to degenerate with those of Kerr BH (Figure 6c), thereby reducing the feasibility of distinguishing between these two types of BHs through gravitational lensing effects. Therefore, from an observational perspective, smaller $k$ values are more suitable for detecting and characterizing slight deviations in BH hair properties.